rank the following fertilizers in decreasing order of mass percentage of nitrogen

KNO3
(NH4)H2PO4
(NH4)2SO4
NH4NO3
NH3
(NHO4)2HPO4
Answer #1

Answer:

Let us calculate Mass % of N in each compound. We will calculate % Mass of N in Mass equal to molar mass of each compound.

Atoomic mass of Nittrogen (N) = 14 g/mole.

1) KNO3 Molar mass = 101.1

% Mass of N = (Mass of N / Molar mass of KNO3) x 100

= (14/101.1) x 100

=13.85 ^%

2) (NH4)H2PO4 , Molar mass = 115.02 g/mole

% Mass of N = (14/115.02) x 100 = 12.17 %

3) (NH4)2SO4, Molar mass = 132.14 g/mole

% mass of N = (28/132.14) x 100 = 21.18 %

4) NH4NO3 Molar mass = 80.04 g/mole and Mass of 2N = 2 x 14 = 28

% Mass of N = (28 /80.04) x 100 = 34.98 %

5) NH3 Molar mass = 17 g/mole

% Mass of N = (14 / 17) x 100 = 82.35 %

6) (NH4)2HPO4 Molar mass= 132.06 g/mole

Mass of 2N = 28

% Mass of N = (28/132.06)x 100 = 21.20 %

Based on % Mass of N calculated we have decreasing order of % N mass as,

NH3 (82.35%) > NH4NO3 (34.98%) > (NH4)2HPO4 (21.20%) > (NH4)2SO4 (21.18%) > KNO3 (13.85) > (NH4)H2PO4 (12.17%)

========================XXXXXXXXXXXXXXX==========================